# Skill-aware Mutual Information Optimisation for Generalisation in Reinforcement Learning

### v1

@@ 4.1 The $log$-$K$ curse of $K$-sample MI estimators
To ensure that the $K$-sample estimator is a tight bound of the ground-truth MI, inspired by Lemma 1, we derive three key insights when learning a context encoder: (1) focus on a ground-truth MI that is smaller than $I(x;y)$ and achievable with a smaller sample size; (2) develop a $K$-sample estimator tighter than $I_{\text{InfoNCE}}$; (3) increasing sample quantity $K$, however, this is usually impractical.
A meta-RL agent learns a context encoder by maximising MI between trajectories $\tau_{c}$ and context embeddings $c$. Driven by insight (1), we introduce Skill-aware Mutual Information (SaMI) in Section 4.2, designed to enhance the zero-shot generalisation of downstream RL tasks. Corresponding to insight (2), we propose Skill-aware Noise Contrastive Estimation (SaNCE) to maximise SaMI with limited samples in Section 4.3. Finally, Section 4.4 demonstrates how to equip any Meta-RL agent with SaNCE in practice.
Figure 3: A $K$-sample estimator $I_{\text{InfoNCE}(c;\pi_{c};\tau_{c})}$, with limited sample size $K$, is a loose lower bound of $I(c;\tau_{c})$ and leads to lower performance embeddings. $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ is a lower ground-truth MI, and $I_{\text{SaNCE}}(c;\pi_{c};\tau_{c})$ is a tighter lower bound of $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$.
@@ 4.2 Skill-aware mutual information: a smaller ground-truth MI
An useful tool in learning a versatile agent is to understand when to explore novel skills or switch between existing skills in multi-task settings. To start with, we define skills:
@@
A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$, abbreviated as $\pi_{c}$. If a skill $\pi_{c}$ is conditioned on a state $s_{t}$, we can sample actions $a_{t}\sim\pi(\cdot|c,s_{t})$. If we sample actions from $\pi_{c}$ at consecutive timesteps, we gain a trajectory $\tau_{c}=\{s_{t},a_{t},r_{t},s_{t+1},...,s_{t+T},a_{t+T},r_{t+T}\}$.
After executing skills, an agent should distinguish tasks through environmental feedback to be aware of underlying environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. In this regard, the context encoder $\psi$ should learn by maximising the MI between context embedding $c$, skills $\pi_{c}$, and environmental feedback (i.e., trajectories $\tau_{c}$). We achieve this learning process by maximising the MI $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$, in which we introduce a variable, skill $\pi_{c}$, into $I(c;\tau_{c})$. Formally, we propose a novel MI optimisation objective for a context encoder, Skill-aware Mutual Information (SaMI), which is defined as:
$I_{\text{SaMI}}(c;\pi_{c};\tau_{c})=\mathbb{E}_{p(c,\pi_{c},\tau_{c})}\left\{\log\ \frac{p(\tau_{c}|c,\pi_{c})}{p(\tau_{c})}\right\}.$ (2)
Although we cannot evaluate $p(c,\pi_{c},\tau_{c})$ directly, we approximate it by Monte-Carlo sampling, using $K$ samples from $p(c,\pi_{c},\tau_{c})$. After introducing the variable $\pi_{c}$, $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ contains less information, i.e., $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$ (see proof in Appendix B). So the context encoder $\psi$ has a lower ground-truth MI $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ to approach. $I_{\text{SaMI}}$ also enables the RL agent to embody diverse skills and switch skills based on the context embedding $c$. In other words, the skill $\pi_{c}$ conditioned on different context embeddings $c$ will exhibit various modes of behaviour, such as the Push skill $\pi_{c_{1}}$ (pushing a cube on the table to the target position) or the Pick&Place skill $\pi_{c_{2}}$.
@@
4.3 Skill-aware noise contrastive estimation: a tighter $K$-sample estimator
Despite InfoNCE’s success as a $K$-sample estimator for approximating MI [Laskin et al., 2020, Eysenbach et al., 2022], its learning efficiency plunges due to limited numerical precision, which is called the $\log$-$K$ curse, i.e., $I_{\text{InfoNCE}}\leq\log K\leq I_{\text{SaMI}}\leq I(c;\tau_{c})$ [Chen et al., 2021] (see proof in Appendix B). When $K\to+\infty$, we can expect $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ [Guo et al., 2022]. However, increasing $K$ is too expensive, especially in complex environments with enormous negative sample space. In response, we propose a novel $K$-sample estimator with a reduced required sample space size $K^{*}$ ($K^{*}\ll+\infty$). First, we define $K^{*}$:
Definition 2 ($\mathbf{K^{*}}$)
@@ 4.3 Skill-aware noise contrastive estimation: a tighter $K$-sample estimator
Figure 5: A practical framework for SaNCE. The context encoder embeds the trajectories into context embeddings, in which the queries are generated by context encoder $\psi$ and the negative/positive embeddings are produced by the momentum context encoder $\psi^{*}$. The encoder is updated with gradients from both the SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and the RL loss $\mathcal{L}_{RL}$. The replay buffer is task-specific.
Note that $K^{*}$ is the maximum batch size that can be sampled in contrastive learning. Therefore, to ensure that $I_{\text{InfoNCE}}$ is a tight bound of $I_{\text{SaMI}}$, we require that $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ when $K\to K^{*}$. Under the definition of $K^{*}$, the replay buffer can be divided according to the different context embeddings $c$ and skills $\pi_{c}$ (i.e., observing context embeddings $c$ and skills $\pi_{c}$). In real-world robotic control tasks, the sample space size significantly increases due to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. Taking the sample space of InfoNCE as an example (Figure 4(a)), in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{c_{1}}$ generated after executing the skill $\pi_{c_{1}}$ in task $e_{1}$, and negative samples are trajectories $\{\tau_{c_{2}},...\}$ from other tasks $\{e_{2},...\}$. The permutations and combinations of $N$ environmental features lead to an exponential growth in task number $|c|$, which in turn results in an increase of sample space $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M$.
We introduce a tight $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), which is used to approximate $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ with a reduced $K^{*}_{\text{SaNCE}}$. For SaNCE, both positive samples $\tau_{c_{1}}^{+}$ and negative samples $\tau_{c_{1}}^{-}$ are sampled from the current tasks $e_{1}$, but are generated by executing positive skills $\pi_{c_{1}}^{+}$ and negative skills $\pi_{c_{1}}^{-}$, respectively. Here, a positive skill is intuitively defined by whether it is optimal for the current task $e$, with a more formal definition provided in Section 4.4. For instance, in a cube-moving task under a large friction setting, the agent executes a skill $\pi_{c}^{+}$ after several iterations of learning, and obtains corresponding trajectories $\tau_{c}^{+}$ where the cubes leave the table surface. This indicates that the skill $\pi_{c}^{+}$ is Pick&Place and other skills $\pi_{c}^{-}$ may include Push or Flip (flipping the cube to the goal position), with corresponding trajectories $\tau_{c}^{-}$ where the cube remains stationary or rolls on the table. Formally, we can optimise the $K$-sample lower bound $I_{\text{SaNCE}}$ to approximate $I_{\text{SaMI}}$:
@@
With a skill-aware context encoder $\psi$, we have $I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq\log K\leq I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$. (see proof in Appendix B)
SaNCE is a plug-and-play module for any other NCE. For example, the negative sample space can be more diverse by combining SaNCE and InfoNCE with $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M$ (Figure 4(c)). A detailed analysis of the sample size of $I_{\text{Sa+InfoNCE}}$ can be found in Appendix C.
4.4 Optimal/Sub-optimal skill-aware sampling strategy
In this section, we further define the positive and negative skills and propose a practical trajectory sampling method. SaNCE requires sample trajectories by ovserving skills $\pi_{c}$. However, we cannot access the prior distribution $p(\pi_{c})$, making it impossible to enumerate all distinct skills $\pi_{c}$. Therefore, in the optimal/sub-optimal skill-aware sampling method, positive skills $\pi_{c}^{+}$ are defined as the optimal skills achieving the highest return $\sum_{i=t}^{t+T}R(s_{i},a_{i})$ in task $e$, while negative skills $\pi_{c}^{-}$ encompass all the other sub-optimal skills with lower returns in task $e$. Thus, we can simply sample the trajectory with the ranked highest return from the replay buffer as the positive sample $\tau_{c}^{+}$, and the one with the lowest return as the negative sample $\tau_{c}^{-}$. The SaNCE loss is then minimised to bring the context embeddings of the highest return trajectories closer in the embedding space while distancing those of negative trajectories.
Note that, at the end of the training, the top-ranked trajectories in the ranked replay buffer correspond to positive samples $\tau_{c}^{+}$ with high returns, and the lower-ranked ones are negative samples $\tau_{c}^{-}$ with low returns. However, before the agent is able to achieve high returns, all trajectories are with low returns. Therefore, our SaNCE loss is a soft version of the $K$-sample SaNCE estimator:
@@ 5 Experiments
We use a three-step process to demonstrate the benefits of SaMI in each environment and answer three questions: (1) Does optimising SaMI lead to increased returns during training and zero-shot generalisation (see Table 5.1 and 2)?; (2) Does SaMI help the RL agents to be versatile and embody multiple skills (see Figure 6)?; (3) Can SaNCE overcome the $\log$-$K$ curse in sample-limited scenarios (see Table 5.1 and 2, and Section 5.4)?
@@ 5.1 Experimental setup
Modified benchmarks with multiple environmental features.  [^1] Our modified benchmarks are open-sourced at https://github.com/uoe-agents/Skill-aware-Panda-gym We demonstrate the efficacy of our method using two benchmarks, Panda-gym [Gallouédec et al., 2021] and MuJoCo [Todorov et al., 2012] (details in Sections 5.2 and 5.3). The benchmarks are modified to be influenced by multiple environmental features simultaneously. Environmental features are sampled at the start of each episode during both training and testing phases. During training, we uniform-randomly select a combination of environmental features from a training set. At test time, we evaluate each algorithm in unseen tasks with different environmental features outside the training range. Generalisation performance is measured in two different regimes: moderate and extreme. The moderate regime draws environmental features from a closer range to the training range compared to the extreme. For all our experiments, we report the mean and standard deviation of the models trained over five seeds in both training and test tasks. Further experimental details are available in Appendix D.
Baselines. In our experiments, we primarily compare InfoNCE with SaNCE to demonstrate the performance improvements brought by SaNCE. Therefore, we compare our method with three prevailing and competitive baselines. First, we consider CCM [Fu et al., 2021], which is equipped with InfoNCE. Additionally, we consider TESAC [Yu et al., 2020], which employs a value function loss, allowing us to evaluate the impact on the context encoder without contrastive loss. Given that CCM and TESAC are using RNN encoder, we also consider PEARL [Rakelly et al., 2019], which utilises an MLP context encoder and a similar loss as TESAC.
Our methods.  [^2] Our code and experimental data are available at https://github.com/uoe-agents/SaMI We employ Soft Actor-Critic (SAC) [Haarnoja et al., 2018] as the base RL algorithm. SaNCE can be integrated with any Meta-RL algorithm. Hence, we equip two RL algorithms with SaNCE: (1) SaTESAC is TESAC with SaNCE, which solely uses SaNCE for contrastive learning, with a significantly smaller sample space than that of other algorithms, as shown in Figure 4(b); (2) SaCCM is CCM with SaNCE, where the contrastive learning combines InfoNCE and SaNCE, as shown in Figure 4(c). The purpose is to compare both the impact of different negative sample spaces on the contrastive learning and downstream RL tasks, and the adaptability of SaNCE across different algorithms. In Appendix G, we provide a comparison in MuJoCo with numerical results from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] because of the exact same environmental setting. The sample spaces of PEARL, TESAC, and CCM are all $|c|$ times larger than that of SaTESAC.

### v3

@@ 4.2 Skill-aware mutual information: a smaller ground-truth MI
We aim for our MI learning objective to incentivise agents to acquire a diverse set of skills, enabling them to generalise effectively across tasks. To start with, we define skills [Eysenbach et al., 2018]:
@@
A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$, abbreviated as $\pi_{c}$. If a skill $\pi_{c}$ is conditioned on a state $s_{t}$, we can sample actions $a_{t}\sim\pi(\cdot|c,s_{t})$. After sampling actions from $\pi_{c}$ at consecutive timesteps, we obtain a trajectory $\tau_{c,t:t_{T}}=\{s_{t},a_{t},r_{t},s_{t+1},\ldots,s_{t+T},a_{t+T},r_{t+T}\}$ which demonstrates a consistent mode of behaviour.
After a limited amount of exploration, an agent should be able to infer the task (i.e., environmental features $e={e^{0},e^{1},\ldots,e^{N}}$) and adapt accordingly within the current episode. The context embedding should encompass skill-related information, guiding the policy on when to explore new skills or switch between existing ones. We propose that the context encoder $\psi$ should be trained by maximising the MI between the context embedding $c$, skills $\pi_{c}$, and trajectories $\tau_{c}$. To this end, we propose a novel MI optimisation objective, Skill-aware Mutual Information (SaMI), defined as:
$I_{\text{SaMI}}(c;\pi_{c};\tau_{c})=I(c;\tau_{c})-I(c;\tau_{c}|\pi_{c}).$ (2)
SaMI is defined according to interaction information [McGill, 1954], serving as a generalisation of MI for three variables $\{c,\pi_{c},\tau_{c}\}$. Although we cannot evaluate $p(c,\pi_{c},\tau_{c})$ directly, we approximate it by Monte-Carlo sampling, using $K$ samples from $p(c,\pi_{c},\tau_{c})$. As illustrated in Figure 3, a context encoder $\psi$ trained with the objective of maximising $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ converges more quickly, as $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$ (see proof in Appendix B). By focusing more on skill-related information, $I_{SaMI}$ enables agents to autonomously discover a diverse range of skills for handling multiple tasks.
@@
4.3 Skill-aware noise contrastive estimation: a tighter $\boldsymbol{K}$-sample estimator
Despite InfoNCE’s success as a $K$-sample estimator for approximating MI [Laskin et al., 2020, Eysenbach et al., 2022], its learning efficiency plunges due to limited numerical precision, which is called the $\log$-$K$ curse, i.e., $I_{\text{InfoNCE}}\leq\log K\leq I_{\text{SaMI}}$ [Chen et al., 2021] (see proof in Appendix B). When $K\to+\infty$, we can expect $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ [Guo et al., 2022]. However, increasing $K$ is too expensive, especially in complex environments with enormous negative sample space. To address this, we propose a novel $K$-sample estimator that requires a significantly smaller sample size $K\ll+\infty$. First, we define $K^{*}$:
Definition 2 ($\boldsymbol{K^{*}}$)
@@ 4.3 Skill-aware noise contrastive estimation: a tighter $\boldsymbol{K}$-sample estimator
Figure 5: A practical framework for using SaNCE in the meta-training phase. During meta-training, we sample trajectories from the replay buffer for off-policy training. Queries are generated by a context encoder $\psi$, which is updated with gradients from both the SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and the RL loss $\mathcal{L}_{RL}$. Negative/Positive embeddings are encoded by a momentum context encoder $\psi^{*}$, which is driven by a momentum update with the encoder $\psi$. During meta-testing, the meta-trained context encoder $\psi$ embeds the current trajectory, and the RL policy takes the embedding as input together with the state for adaptation within an episode.
To ensure that $I_{\text{InfoNCE}}$ is a tight bound of $I_{\text{SaMI}}$, we require that $I_{\text{InfoNCE}}\approx\log K\approx I_{\text{SaMI}}$ when $K\to K^{*}$. Under the definition of $K^{*}$, the replay buffer can be divided according to the different context embeddings $c$ and skills $\pi_{c}$ (i.e., observing context embeddings $c$ and skills $\pi_{c}$). In real-world robotic control tasks, the sample space size significantly increases due to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. Taking the sample space of InfoNCE as an example (Figure 4(a)), in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{c_{1}}$ generated after executing the skill $\pi_{c_{1}}$ in task $e_{1}$, and negative samples are trajectories $\{\tau_{c_{2}},...\}$ from other tasks $\{e_{2},...\}$. The permutations and combinations of $N$ environmental features lead to an exponential growth in task number $|c|$, which in turn results in an increase of sample space $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M$.
We introduce a tight $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), which is used to approximate $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ with $K^{*}_{\text{SaNCE}}<K^{*}_{\text{InfoNCE}}$. For SaNCE, both positive samples $\tau_{c_{1}}^{+}$ and negative samples $\tau_{c_{1}}^{-}$ are sampled from the current tasks $e_{1}$, but are generated by executing positive skills $\pi_{c_{1}}^{+}$ and negative skills $\pi_{c_{1}}^{-}$, respectively. Here, a positive skill is intuitively defined by whether it is optimal for the current task $e$, with a more formal definition provided in Section 4.4. For instance, in a cube-moving task under a large friction setting, the agent executes a skill $\pi_{c}^{+}$ after several iterations of learning, and obtains corresponding trajectories $\tau_{c}^{+}$ where the cubes leave the table surface. This indicates that the skill $\pi_{c}^{+}$ is Pick&Place and other skills $\pi_{c}^{-}$ may include Push or Flip (flipping the cube to the goal position), with corresponding trajectories $\tau_{c}^{-}$ where the cube remains stationary or rolls on the table. Formally, we can optimise the $K$-sample lower bound $I_{\text{SaNCE}}$ to approximate $I_{\text{SaMI}}$:
@@
With a context encoder $\psi$ and finite sample size $K$, we have $I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)\leq\log K\leq I_{\text{SaMI}}(c;\pi_{c};\tau_{c})\leq I(c;\tau_{c})$. (see proof in Appendix B)
SaNCE can be used alone or combined with other optimisation objectives to train context encoders in Meta-RL algorithms. For instance, integrating SaNCE with InfoNCE diversifies the negative sample space, with $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M$. The sample space for $I_{\text{Sa+InfoNCE}}$ is depicted in Figure 4(c) and further analysed in detail in Appendix C.
4.4 Skill-aware trajectory sampling strategy
In this section, we propose a practical trajectory sampling method. Methods focusing on skill diversity often rely heavily on accurately defining and identifying individual skills [Eysenbach et al., 2018]. Some of these methods require a prior skill distribution, which is often inaccessible [Shi et al., 2022], and it is impractical to enumerate all possible skills that we hope the model to learn. Besides, we believe that distinctiveness of skills is inherently difficult to achieve — a slight difference in states can make two skills distinguishable, and not necessarily in a semantically meaningful way. Consequently, we do not directly teach any of these skills or assume any prior skill distribution. Diverse skills naturally emerge from the incentives of the SaMI learning objective in a multi-task setting, driven by the inherent need to develop generalisable skills. For example, in high-friction tasks, the agent must acquire the Pick&Place skill to avoid large frictional forces, whereas in high-mass tasks, the agent must learn the Push skill since it cannot lift the cube. In each task, we only identify whether the skills are optimal; in this way, under a multi-task setting, the agent will acquire a set of general skills that are applicable to many tasks.
In a given task $e$, positive skills $\pi_{c}^{+}$ are defined as optimal skills achieving highest return $\sum_{i=t}^{t+T}R(s_{i},a_{i})$, whereas negative skills $\pi_{c}^{-}$ are those that result in lower returns. As a result, the positive sample $\tau_{c}^{+}$ consists of trajectories generated by the skills with the highest ranked returns, while the negative samples correspond to those with the lowest returns. This straightforward approach of selecting positive samples based on the ranked highest return effectively aligns with positive skills and mitigates the challenge of hard negative examples [Robinson et al., 2021]. The SaNCE loss is then minimised to bring the context embeddings of the highest return trajectories closer while distancing those of negative trajectories. By the end of training, the top-ranked trajectories in the ranked replay buffer correspond to positive samples $\tau_{c}^{+}$ with high returns, while the lower-ranked trajectories represent negative samples $\tau_{c}^{-}$ with low returns. However, at the beginning of training, it is likely that all trajectories have low returns. Therefore, our SaNCE loss is a soft variant of the $K$-sample SaNCE:
@@ 5 Experiments
Our experiments aim to answer the following questions: (1) Does optimising SaMI lead to increased returns during training and zero-shot generalisation (see Table 5.1 and 2)?; (2) Does SaMI help the RL agents to be versatile and embody multiple skills (see Figure 6)?; (3) Can SaNCE overcome the $\log$-$K$ curse in sample-limited scenarios (see Table 5.1 and 2, and Section 5.4)?
Figure 6: (a) UMAP visualisation of context embeddings for the SaCCM in the Panda-gym environment, with points in the yellow box representing the Push skill in high-mass tasks. Heatmap of (b) success rate, (c) Push skill probability, and (d) Pick&Place skill probability for SaCCM. In large-mass scenarios, the Push skill is more likely to be executed than Pick&Place.
@@ 5.1 Experimental setup
Modified benchmarks with multiple environmental features. [^1] Our modified benchmarks are open-sourced at https://github.com/uoe-agents/Skill-aware-Panda-gym We evaluate our method on two benchmarks, Panda-gym [Gallouédec et al., 2021] and MuJoCo [Todorov et al., 2012] (details in Sections 5.2 and 5.3). The benchmarks are modified to be influenced by multiple environmental features, which are sampled at the start of each episode during meta-training and meta-testing. During meta-training, we uniform-randomly select a combination of environmental features from a training task set. At test time, we evaluate each algorithm in unseen tasks with environmental features outside the training range. Generalisation performance is measured in two different regimes: moderate and extreme. The moderate regime draws environmental features from a closer range to the training range compared to the extreme. Our results report the mean and standard deviation of models trained over five seeds in both training and test tasks. Further details are available in Appendix D.
Baselines. The loss function of the context encoder in all algorithms consists of two key components: the RL loss and the contrastive loss. The RL loss $\mathcal{L}_{RL}$, which is the same across all methods, corresponds to the RL value function loss. Our primary comparison focuses on the contrastive loss, taking the form of either SaNCE, InfoNCE, or no contrastive loss. Accordingly, we select baselines based on contrastive loss: CCM [Fu et al., 2021], which utilises InfoNCE, and TESAC [Yu et al., 2020], which relies solely on the RL loss, allowing assessment of the context encoder without contrastive loss. Since CCM and TESAC use an RNN encoder, we also include PEARL [Rakelly et al., 2019], which utilises an MLP context encoder and follows the similar RL loss. Additionally, Appendix G includes comparisons with DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020], using the same environmental setup in the MuJoCo benchmark.
Our method. [^2] Our code, video demos and experimental data are available at https://github.com/uoe-agents/SaMI We use Soft Actor-Critic (SAC) [Haarnoja et al., 2018] as the base RL algorithm, training agents for 1.6 million timesteps in each environment (details in Appendix D.3). SaNCE is a simple objective based on MI that can be used to train any context encoder. We integrate SaNCE into two Meta-RL algorithms: (1) SaTESAC is TESAC with SaNCE, which uses SaNCE for contrastive learning, using a $|c|$ times smaller sample space (Figure 4(b)); (2) SaCCM is CCM with SaNCE, where the contrastive learning combines InfoNCE and SaNCE, as shown in Figure 4(c).